Consider the table. Fixed Cost ($) 10,000 10,000 10,000 10,000 10,000 10,000 10,000 Labor Quantity Variable Cost ($)Total Cost (S)Marginal Cost ($)Average Cost ($)AVC* ($) (hours) 10,000 15,000 40 45 65 78 5,000 10,000 15,000 277.78 222.22 230.77 256.41 284.09 50 185.19 444.44 25,000 30,000 35,000 40,000 250.00 384.62 70 80 90 25,000 30,000 397.73 425.53 500.00 94. *Average Variable Cost The table shows the cost structure of a firm producing computer mainframes Calculate the missing values A through H and enter these into the boxes provided
Answer #1

ANSWER:

A) Variable cost = total cost - fixed cost = 30,000 - 10,000 = 20,000

B) Total cost = fixed cost + variable cost = 10,000 + 10,000 = 20,000

C) Marginal cost = change in total cost / change in quantity = (15,000 - 10,000) / (18 - 0) = 5,000 / 18 = 277.77

D) Marginal cost = change in total cost / change in quantity = (30,000 - 25,000) / (78 - 65) = 5,000 / 13 = 384.61

E) Marginal cost = change in total cost / change in quantity = (40,000 - 35,000) / (94 - 88) = 5,000 / 6 = 833.33

F) Average cost = Total cost / quantity = 15,000 / 18 = 833.33

G) Average cost = total cost / quantity = 30,000 / 78 = 384.61

H) Average variable cost = variable cost / quantity = 30,000 / 94 = 319.14